# Neural Networks for Fast Optimisation in Model Predictive Control: A Review

### v2

@@ 5.2 Methods with probabilistic guarantees
In this section, we discuss a method for generating NNBO approximations of NMPC with probabilistic guarantees on constraint satisfaction based on Hoeffding’s Inequality (HI). We start by providing a brief explanation of this inequality because other studies discussed hereafter are also based on it. HI, as shown in Equation (6), defines the upper bound of the probability of the mean of a set of $n$ randomly sampled variables $(\xi_{i})$ from a distribution being further than a value $(\epsilon)$ from the mean of the distribution itself $(E(\xi))$ Von Luxburg and Schölkopf (2011). The value of the upper bound is given by the RHS of Equation (6). Thus note that for a fixed $\epsilon$ the upper bound decreases with the number of samples. This statistical result can be applied in several ways when generating an NNBO approximation of MPC to yield different kinds of guarantees. For instance, it is possible to train DNN optimisers that find solutions with bounded levels of suboptimality with a certain probability, or DNN optimisers that guarantee constraint satisfaction with some probability, or DNN optimisers that guarantee stability.
@@
Cao and Gopaluni Cao and Gopaluni (2020) and Li et al. Li et al. (2022) used HI to come up with a framework for generating NNBO approximations of NMPC with probabilistic guarantees on constraint satisfaction. Similar to methods with MPC-inspired loss functions, their method embeds the MPC optimisation problem into the DNN training problem and in this case it results in a unified nonlinear two-stage stochastic optimisation problem. To provide the statistical guarantees, Cao and Gopaluni Cao and Gopaluni (2020) used the same statistical learning theory principles as Zhang et al. Zhang et al. (2020), but instead of applying it to point samples they applied it to closed-loop trajectory samples. That idea was first proposed in the work by Hertneck et al Hertneck et al. (2018) which will be discussed in Section 6.2. In this case an appropriate choice of number of samples leads to a probabilistic guarantee of constraint satisfaction for the DNN approximation for any trajectory starting from any point in the prescribed distribution of starting points. Similar to Zhang et al.’s Zhang et al. (2020) approach, the calculated number of required samples is usually quite large. Therefore, Li et al. Li et al. (2022) employed a heuristic based on HI where DNN approximations for an increasing number of sample trajectories are produced until satisfactory performance and guarantees are achieved. In this study two methods were proposed to solve the “optimise and train” nonlinear two-stage stochastic optimisation problem, namely, a numeric interior point based solver, and an equivalent RNN training problem. Both methods were demonstrated on a water tank level control problem with nonlinear dynamics, four states, two outputs, and two inputs. The resulting DNN approximation sped up optimisation by three orders of magnitude compared to a conventional solver. Li et al. Li et al. (2022) also showed how the performance of the approximation improves with the incorporation of additional trajectories and carefully selected critical scenarios during the training process. Importantly, as more samples are added, the interior point solver eventually becomes unable to solve the problem. As a result, for large data sets only the RNN optimisation approach is viable. The sole shortcoming of the study is not considering more than one benchmark problem to evaluate the scalability and robustness of the approach.
@@ 5.3 Methods with strict guarantees
In this section, we cover methods and frameworks that can be used to generate NNBO approximations of NMPC in ways that strictly preserve some or all of the desirable properties of the original controller such as constraint satisfaction, stability and recursive feasibility. Following their pioneering work Parisini and Zoppoli (1995) on learning MPC policies with NNs, Parisini et al. continued to develop their ideas and eventually addressed the problem of constrained NMPC. In Parisini et al. (1998), they proposed a set of optimisation routines and heuristic iterative procedures that could be used to compute bounds on the terminal weights and number of neurons required to create NMPC approximations with bounded output error over the set of states for which the approximation was generated. With sufficiently low bounds on the output error the approximate NMPC controller is guaranteed to stabilise the system to a small user-prescribed region around the origin, thus making this approach the first with guarantees on stability. The proposed methods were demonstrated on a small plant. However, on a second example involving a larger and more complex plant, the authors omitted all of the procedures that yield the stability guarantees and went back to the pure regression approach, presumably because the problems involved were intractable for large systems. This line of work was continued over a decade later by Pin et al. in  Pin et al. (2010) and  Pin et al. (2013), where the stability properties of the approximate control laws were further studied in the presence of approximation error and modelling uncertainties. As reported by Parisini et al. Parisini et al. (1998), these studies showed that the approximate control policies generated with this framework are only able to stabilise systems to a region near the origin. Furthermore, no new evidence on the scalability of the approaches was provided. Instead, both studies used the same example system, a nonlinear oscillator, featuring a small plant with two states and one input. The prediction and control horizons considered in the exact NMPC controllers were also short, as outlined in Table 2. In fact, Hertneck et al Hertneck et al. (2018) reported that for the framework in Pin et al. (2013) the admissible approximation error required to yield guarantees was typically unachievable.
@@
In the same year when Parisini et al. Parisini et al. (1998) formalised their framework, Ahmed and Dajani Ahmed and Al-Dajani (1998) proposed a fundamentally different approach, where instead of learning the NMPC policy via regression, the loss function of the approximating NN was set to the cost function of a regulating NMPC controller. The resulting NN could be trained using back propagation by calculating the required derivatives using the concept of block partial derivatives. The authors also introduced an important constraint to ensure stability at the origin. Specifically, they set out to train a network that would produce no output if the system was at the equilibrium position. As they point out, a linear control law with no offset achieves this by design, so the architecture of the approximating NN was set to a deep NN with no bias terms and $tanh$ activation such that if all inputs to the network were zero, the outputs would also be zero. The proposed approach was shown to work for state and output regulation tasks of small linear and nonlinear plants with prediction and control horizons of up to 300 steps, with and without input disturbance. While those simulation results are very promising, it should be noted that the approach has some disadvantages. For instance, training requires human supervision at every step to ensure stability. This involves manually adjusting the network architecture and learning rate as well as gradually increasing the length of the prediction horizon. Also, the only guarantees that the approach Ahmed and Al-Dajani (1998) provides are input constraint satisfaction via saturation and stability at the equilibrium point. State and output constraint satisfaction are not guaranteed and neither is stability during regulation.
@@
Some works have focused on the approximation of NMPC controllers with binary control inputs. Lohr et al. Löhr et al. (2020) considered the problem of mixed-integer MPC with binary decision variables for linear state space plant models. Despite considering a linear state space plant, the use of binary control variables makes this problem nonlinear. Typically, this type of MPC would require solving a Mixed Integer Quadratic Program (MIQP) which is much slower than solving a standard QP. To get around this, Lohr et al. proposed training a NN to predict the value of all binary variables at the first step in the prediction horizon. Moreover, they relaxed the binary constraint for all other steps in the prediction horizon by allowing the binary variables to be continuous but limited to the range [0, 1]01[0,\;1][ 0 , 1 ]. This effectively converted the MIQP into a standard QP that could be solved much quicker. Given that the resulting approximate strategy still involves solving a QP in real time, this approximation framework preserves all the theoretical properties of LMPC. The effectiveness of the approach Löhr et al. (2020) was demonstrated on a domestic heating control problem with a small plant and two binary control inputs. For that problem, the performance of the approximate controller was on par with that of the original controller where the MIQP was solved using a conventional solver. Nonetheless, the approximate controller ran three times faster.
@@
Vaupel et al. Vaupel et al. (2020) further explored the concept of using NN approximations alongside conventional optimisation methods. They studied two strategies: initialising the NMPC solution from the outputs obtained from a DNN approximation of the controller, and refining the solution of DNN approximations by solving a QP that approximates the original NMPC problem. Both of these approaches require training of a DNN approximation of the original NMPC controller. In this case the approximation needs to predict the value of all control actions at each step along the control horizon of the controller. To achieve this the authors used an imitation learning setup with an MSE loss function and training data generated from running closed loop simulations of the NMPC controller. The inputs of the network were set to the states, references, and disturbances at the current control interval. In the NMPC initialisation case, the control actions produced by the DNN approximation were used to hot start a conventional optimisation solver which went on to solve the NMPC problem to convergence. In the second case, the outputs of the DNN were used to construct a QP that was an approximation of the full NLP problem. This approximation was generated by computing the hessian of the Lagrangian and gradients of the objective and constraint functions for the control action values predicted by the DNN approximation. The QP was designed such that its solution satisfied a first order approximation of the constraints of the problem thereby strictly preserving the theoretical guarantees on MPC. In practice guaranteeing the convexity and feasibility of the resulting QP requires additional strategies such as modifying the calculated Hessian until it is positive definite or solving a feasibility restoring QP. In both cases the output of the post processing QP is a set of deltas that can be added to the original guess of the DNN approximation to improve it and make it constraint compliant. Through simulation studies on a small MIMO plant and an original NMPC controller with short prediction and control horizons, Vaupel et al. Vaupel et al. (2020) reported that using the DNN outputs to initialise the NMPC solution slowed down the controller by about 45% on average. However, the maximum solution time decreased by 42% and computation times were much more consistent. Thus, while this approach could be useful for some applications, it is not relevant for the purpose of significantly speeding up existing controllers. On the other hand, the post processing QP approach was on average 24 times faster than solving the full NMPC problem and it led to negligible to no deterioration in control performance. When compared against the popular approach of generating an NMPC approximation via pure regression, the resulting strategy was able to respect all control constraints and maintain close to optimal cost values during operation. In contrast, the simple regression approach violated constraints and exhibited running costs far from optimal.
@@
Bonzanini et al. Bonzanini et al. (2020) proposed another post processing method relying on projection operations instead of further optimisation. In this method the outputs of a DNN approximation of NMPC are projected onto the Robust Admissible Input (RAI) set. This projection ensures adherence to all input constraints and guarantees that system states remain within the Robust Control Invariant (RCI) set, which is state constraint compliant by design. The projection operation consists of solving a secondary optimisation problem at every control interval, but said optimisation problem can be formulated as a standard QP and solved offline as a multi-parametric QP (mpQP). The problem with this approach, as with other of similar nature, is that mpQP does not scale well to large problems and for large cases the lookup operation itself can be infeasible in real time at high control rates, especially for embedded systems where computational resources are limited. Indeed the example provided by Bonzanini et al. Bonzanini et al. (2020) was for a small plant with only 2 states and 2 inputs, but it should be noted that the length of the prediction and control horizons was quite significant at 100 steps. Thus, while the scalability of the method with plant size is unknown, for small plants the method scales well to long horizons.
@@
Instead of post processing the output of a DNN approximation, Wang et al. Wang et al. (2022) proposed an imitation learning setup where the resulting network respects the constraints of the original NMPC problem by design. This result was achieved by setting the loss function of the DNN to a function corresponding to the dual of a problem which minimises the difference in cost between the approximate and original controllers subject to the constraints of the original controller. The training problem then becomes a max-min problem where the goal is to maximise the dual variables while minimising the loss by adjusting the parameters of the DNN. Wang et al. solved this problem using dual gradient descent and a heuristic algorithm that progressively expanded the size of the training data set until a desired level of performance was achieved which guaranteed input and state constraint compliance of the approximation.
@@ 6 Neural Network Based Optimisation for Robust Model Predictive Control
Lucia and Karg (2018) 1a
Train a DNN via simple imitation learning with MSE loss to learn scenario RMPC
None
7 1 3 15 15 1 2,800x Sim
Karg and Lucia (2019) 1e
Train DNN to imitate msNMPC, use constraint back-off parameters during training data generation to improve likelihood of constraint satisfaction.
None
3 3 1 80 80 1 Sim
Bonzanini et al. (2021) 1a
Use GPs to learn plant model time-varying mismatch and adapt scenario tree of scenario MPC online. Train DNN with MSE loss to imitate resulting controller.
None
3 1 2 5 5 1 4.4x Sim
Karg and Lucia (2021) 3c
Train a DNN with simple MSE loss to imitate real MPC controller, use the network to initialise an RL agent, improve RL agent via further training.
None
7 1 3 20 20 1 Sim
Hertneck et al. (2018) 2b
Apply HI to performance of DNN approximation over trajectories to ensure stability with high probability.
Probabilistic stability and constraint satisfaction based on HI
2 2 1 180 180 200x Sim
Nubert et al. (2020) 2b
Apply HI to performance of DNN approximation over trajectories to ensure stability with high probability.
Probabilistic stability and constraint satisfaction based on HI
8 3 4 200x Robot
Karg et al. (2021) 2b
Finite families based method for obtaining probabilistic guarantees of controllers using non-binary performance indicators.
Probabilistic constraint satisfaction
3 1 1 40 40 1 Sim
Daosud et al. (2019) 3a
Train DNN to predict cost of msRMPC, use this prediction as terminal cost of reformulated controller with shorter horizon.
Stability and constraint satisfaction but not optimality
7 4 2 15 15 1 2.3x Sim
Paulson and Mesbah (2020) 3b
Use a projection operation to coerce the outputs of a DNN approximation into sets that provide theoretical guarantees. Formulate the projection operation as a pQP that can be solved offline.
Constraint satisfaction and under special conditions, input-to-state stability.
2 1 1 10 10 2 400x Sim

### v3

@@ 5.2 Methods with probabilistic guarantees
In this section, we discuss a method for generating NNBO approximations of NMPC with probabilistic guarantees on constraint satisfaction based on Hoeffding’s Inequality (HI). We start by providing a brief explanation of this inequality because other studies discussed hereafter are also based on it. HI, as shown in (6), defines the upper bound of the probability of the mean of a set of $n$ randomly sampled variables $(\xi_{i})$ from a distribution being further than a value $(\epsilon)$ from the mean of the distribution itself $(E(\xi))$ (Von Luxburg and Schölkopf, 2011). The value of the upper bound is given by the RHS of (6). Thus note that for a fixed $\epsilon$ the upper bound decreases with the number of samples. This statistical result can be applied in several ways when generating an NNBO approximation of MPC to yield different kinds of guarantees. For instance, it is possible to train DNN optimisers that find solutions with bounded levels of suboptimality with a certain probability, or DNN optimisers that guarantee constraint satisfaction with some probability, or DNN optimisers that guarantee stability.
@@
Cao and Gopaluni (2020) and Li et al. (2022) used HI to come up with a framework for generating NNBO approximations of NMPC with probabilistic guarantees on constraint satisfaction. Similar to methods with MPC-inspired loss functions, their method embeds the MPC optimisation problem into the DNN training problem and in this case it results in a unified nonlinear two-stage stochastic optimisation problem. To provide the statistical guarantees, Cao and Gopaluni (2020) used the same statistical learning theory principles as Zhang et al. (2020), but instead of applying it to point samples they applied it to closed-loop trajectory samples. That idea was first proposed in the work by Hertneck et al. (2018) which will be discussed in Section 6.2. In this case an appropriate choice of number of samples leads to a probabilistic guarantee of constraint satisfaction for the DNN approximation for any trajectory starting from any point in the prescribed distribution of starting points. Similar to Zhang et al. (2020), the calculated number of required samples is usually quite large. Therefore, Li et al. (2022) employed a heuristic based on HI where DNN approximations for an increasing number of sample trajectories are produced until satisfactory performance and guarantees are achieved. In this study two methods were proposed to solve the “optimise and train” nonlinear two-stage stochastic optimisation problem, namely, a numeric interior point based solver, and an equivalent RNN training problem. Both methods were demonstrated on a water tank level control problem with nonlinear dynamics, four states, two outputs, and two inputs. The resulting DNN approximation sped up optimisation by three orders of magnitude compared to a conventional solver. Li et al. (2022) also showed how the performance of the approximation improves with the incorporation of additional trajectories and carefully selected critical scenarios during the training process. Importantly, as more samples are added, the interior point solver eventually becomes unable to solve the problem. As a result, for large data sets only the RNN optimisation approach is viable. The sole shortcoming of the study is not considering more than one benchmark problem to evaluate the scalability and robustness of the approach.
@@ 5.3 Methods with strict guarantees
In this section, we cover methods and frameworks that can be used to generate NNBO approximations of NMPC in ways that strictly preserve some or all of the desirable properties of the original controller such as constraint satisfaction, stability and recursive feasibility. Following their pioneering work (Parisini and Zoppoli, 1995) on learning MPC policies with NNs, Parisini et al. continued to develop their ideas and eventually addressed the problem of constrained NMPC. In Parisini et al. (1998), they proposed a set of optimisation routines and heuristic iterative procedures that could be used to compute bounds on the terminal weights and number of neurons required to create NMPC approximations with bounded output error over the set of states for which the approximation was generated. With sufficiently low bounds on the output error the approximate NMPC controller is guaranteed to stabilise the system to a small user-prescribed region around the origin, thus making this approach the first with guarantees on stability. The proposed methods were demonstrated on a small plant. However, on a second example involving a larger and more complex plant, the authors omitted all of the procedures that yield the stability guarantees and went back to the pure regression approach, presumably because the problems involved were intractable for large systems. This line of work was continued over a decade later by Pin et al. (2010, 2013), where the stability properties of the approximate control laws were further studied in the presence of approximation error and modelling uncertainties. As reported by Parisini et al. (1998), these studies showed that the approximate control policies generated with this framework are only able to stabilise systems to a region near the origin. Furthermore, no new evidence on the scalability of the approaches was provided. Instead, both studies used the same example system, a nonlinear oscillator, featuring a small plant with two states and one input. The prediction and control horizons considered in the exact NMPC controllers were also short, as outlined in Table 2. In fact, Hertneck et al. (2018) reported that for the framework in Pin et al. (2013) the admissible approximation error required to yield guarantees was typically unachievable.
@@
In the same year when Parisini et al. (1998) formalised their framework, Ahmed and Al-Dajani (1998) proposed a fundamentally different approach, where instead of learning the NMPC policy via regression, the loss function of the approximating NN was set to the cost function of a regulating NMPC controller. The resulting NN could be trained using back propagation by calculating the required derivatives using the concept of block partial derivatives. The authors also introduced an important constraint to ensure stability at the origin. Specifically, they set out to train a network that would produce no output if the system was at the equilibrium position. As they point out, a linear control law with no offset achieves this by design, so the architecture of the approximating NN was set to a deep NN with no bias terms and $tanh$ activation such that if all inputs to the network were zero, the outputs would also be zero. The proposed approach was shown to work for state and output regulation tasks of small linear and nonlinear plants with prediction and control horizons of up to 300 steps, with and without input disturbance. While those simulation results are very promising, it should be noted that the approach has some disadvantages. For instance, training requires human supervision at every step to ensure stability. This involves manually adjusting the network architecture and learning rate as well as gradually increasing the length of the prediction horizon. Also, the only guarantees that the approach by Ahmed and Al-Dajani (1998) provides are input constraint satisfaction via saturation and stability at the equilibrium point. State and output constraint satisfaction are not guaranteed and neither is stability during regulation.
@@
Some works have focused on the approximation of NMPC controllers with binary control inputs. Löhr et al. (2020) considered the problem of mixed-integer MPC with binary decision variables for linear state space plant models. Despite considering a linear state space plant, the use of binary control variables makes this problem nonlinear. Typically, this type of MPC would require solving a Mixed Integer Quadratic Program (MIQP) which is much slower than solving a standard QP. To get around this, the authors proposed training a NN to predict the value of all binary variables at the first step in the prediction horizon. Moreover, they relaxed the binary constraint for all other steps in the prediction horizon by allowing the binary variables to be continuous but limited to the range [0, 1]01[0,\;1][ 0 , 1 ]. This effectively converted the MIQP into a standard QP that could be solved much quicker. Given that the resulting approximate strategy still involves solving a QP in real time, this approximation framework preserves all the theoretical properties of LMPC. Löhr et al. (2020) demonstrated the effectiveness of the approach on a domestic heating control problem with a small plant and two binary control inputs. For that problem, the performance of the approximate controller was on par with that of the original controller where the MIQP was solved using a conventional solver. Nonetheless, the approximate controller ran three times faster.
@@
Vaupel et al. (2020) further explored the concept of using NN approximations alongside conventional optimisation methods. They studied two strategies: initialising the NMPC solution from the outputs obtained from a DNN approximation of the controller, and refining the solution of DNN approximations by solving a QP that approximates the original NMPC problem. Both of these approaches require training of a DNN approximation of the original NMPC controller. In this case the approximation needs to predict the value of all control actions at each step along the control horizon of the controller. To achieve this the authors used an imitation learning setup with an MSE loss function and training data generated from running closed loop simulations of the NMPC controller. The inputs of the network were set to the states, references, and disturbances at the current control interval. In the NMPC initialisation case, the control actions produced by the DNN approximation were used to hot start a conventional optimisation solver which went on to solve the NMPC problem to convergence. In the second case, the outputs of the DNN were used to construct a QP that was an approximation of the full NLP problem. This approximation was generated by computing the hessian of the Lagrangian and gradients of the objective and constraint functions for the control action values predicted by the DNN approximation. The QP was designed such that its solution satisfied a first order approximation of the constraints of the problem thereby strictly preserving the theoretical guarantees on MPC. In practice guaranteeing the convexity and feasibility of the resulting QP requires additional strategies such as modifying the calculated Hessian until it is positive definite or solving a feasibility restoring QP. In both cases the output of the post processing QP is a set of deltas that can be added to the original guess of the DNN approximation to improve it and make it constraint compliant. Through simulation studies on a small MIMO plant and an original NMPC controller with short prediction and control horizons, Vaupel et al. (2020) reported that using the DNN outputs to initialise the NMPC solution slowed down the controller by about 45% on average. However, the maximum solution time decreased by 42% and computation times were much more consistent. Thus, while this approach could be useful for some applications, it is not relevant for the purpose of significantly speeding up existing controllers. On the other hand, the post processing QP approach was on average 24 times faster than solving the full NMPC problem and it led to negligible to no deterioration in control performance. When compared against the popular approach of generating an NMPC approximation via pure regression, the resulting strategy was able to respect all control constraints and maintain close to optimal cost values during operation. In contrast, the simple regression approach violated constraints and exhibited running costs far from optimal.
@@
Bonzanini et al. (2020) proposed another post processing method relying on projection operations instead of further optimisation. In this method the outputs of a DNN approximation of NMPC are projected onto the Robust Admissible Input (RAI) set. This projection ensures adherence to all input constraints and guarantees that system states remain within the Robust Control Invariant (RCI) set, which is state constraint compliant by design. The projection operation consists of solving a secondary optimisation problem at every control interval, but said optimisation problem can be formulated as a standard QP and solved offline as a multi-parametric QP (mpQP). The problem with this approach, as with other of similar nature, is that mpQP does not scale well to large problems and for large cases the lookup operation itself can be infeasible in real time at high control rates, especially for embedded systems where computational resources are limited. Indeed, the example provided by Bonzanini et al. (2020) was for a small plant with only 2 states and 2 inputs, but it should be noted that the length of the prediction and control horizons was quite significant at 100 steps. Thus, while the scalability of the method with plant size is unknown, for small plants the method scales well to long horizons.
@@
Instead of post-processing the output of a DNN approximation, Wang et al. (2022) proposed an imitation learning setup where the resulting network respects the constraints of the original NMPC problem by design. This result was achieved by setting the loss function of the DNN to a function corresponding to the dual of a problem which minimises the difference in cost between the approximate and original controllers subject to the constraints of the original controller. The training problem then becomes a max-min problem where the goal is to maximise the dual variables while minimising the loss by adjusting the parameters of the DNN. Wang et al. solved this problem using dual gradient descent and a heuristic algorithm that progressively expanded the size of the training data set until a desired level of performance was achieved which guaranteed input and state constraint compliance of the approximation.
@@ 6 Neural Network Based Optimisation for Robust Model Predictive Control
Lucia and Karg (2018) 1a Train a DNN via simple imitation learning with MSE loss to learn scenario RMPC None 7 1 3 15 15 1 2,800x Sim
Karg and Lucia (2019) 1e Train DNN to imitate msNMPC, use constraint back-off parameters during training data generation to improve likelihood of constraint satisfaction. None 3 3 1 80 80 1 Sim
Bonzanini et al. (2021) 1a Use GPs to learn plant model time-varying mismatch and adapt scenario tree of scenario MPC online. Train DNN with MSE loss to imitate resulting controller. None 3 1 2 5 5 1 4.4x Sim
Karg and Lucia (2021) 3c Train a DNN with simple MSE loss to imitate real MPC controller, use the network to initialise an RL agent, improve RL agent via further training. None 7 1 3 20 20 1 Sim
Hertneck et al. (2018) 2b Apply HI to performance of DNN approximation over trajectories to ensure stability with high probability. Probabilistic stability and constraint satisfaction based on HI 2 2 1 180 180 200x Sim
Nubert et al. (2020) 2b Apply HI to performance of DNN approximation over trajectories to ensure stability with high probability. Probabilistic stability and constraint satisfaction based on HI 8 3 4 200x Robot
Karg et al. (2021) 2b Finite families based method for obtaining probabilistic guarantees of controllers using non-binary performance indicators. Probabilistic constraint satisfaction 3 1 1 40 40 1 Sim
Daosud et al. (2019) 3a Train DNN to predict cost of msRMPC, use this prediction as terminal cost of reformulated controller with shorter horizon. Stability and constraint satisfaction but not optimality 7 4 2 15 15 1 2.3x Sim
Paulson and Mesbah (2020) 3b Use a projection operation to coerce the outputs of a DNN approximation into sets that provide theoretical guarantees. Formulate the projection operation as a pQP that can be solved offline. Constraint satisfaction and under special conditions, input-to-state stability. 2 1 1 10 10 2 400x Sim